OpenStudy (anonymous):

How do you change a repeating decimal to a fraction for example 1.0101.. please? Help!!

OpenStudy (agent0smith):

let n = 1.010101... Multiply it by 100: 100n = 101.010101... now subtract n from that 100n - n = 101.010101... - 1.010101... = 100 100n-n is 99n, so 99n = 100 now divide both sides by 99.